In a certain week's Lotto draw, 4 people each won R250 000; 834 each won R479.30; 17 134 each won R14.30; and 21 804 each won R9.20. If there were 460 224 entrants who won nothing, find the mean and median prize payout.
Answer #1

Solution:

No. Of people Money won

4 250000

834   479.30

17134 14.30

21804 9.20

460224 0

Mean of this data can be calculated as

Mean = summation(No. Of people *money won)/ summation (no.of people) = ((4*250000)+(834*479.30)+(17134*14.30)+(21804*9.20)+(0*460224))/(4+834+17134+21804+460224) = 1845349/500000 = 3.69

Median value = (500000/2) = 250000th value

So median = 0